# A Large-scale Empirical Study on Improving the Fairness of Deep Learning Models

### 08 Jan 2024 v1

@@ 4.1. Dataset Selection
As a result, we selected three datasets (i.e., CelebA, UTKFace, and CIFAR-10S) for conducting the empirical study. The summary of the datasets is shown in Table 2, which presents the sensitive attribute considered in the study, the labels of classification tasks, and the studied methods that previously used the same setting. In addition, we also presents the number of instances for model training and testing per each dataset by following existing studies.
@@
Specifically, CelebA and UTKFace represent human-centric datasets that involve sensitive attributes (e.g., age and race) reflected in facial images, while CIFAR-10S is a more general dataset in fairness study, which introduces biases related to image colors of objects (non-facial images). They were previously used for evaluation by partial/all methods under different conditions, e.g, using different metrics. From the table we can also find that previous studies tend to employ different datasets, hindering their comprehensive comparison. In this study, we aim to provide an extensive study through a systematic evaluation by employing all of the them, which vary in multiple perspectives.
@@ 4.2.1. Fairness Metrics
Numerous prior studies have predominantly assessed fairness using a restricted set of fairness metrics. Although certain approaches have introduced their unique fairness metrics, these metrics have not achieved widespread adoption and lack broad applicability. To address these limitations, we employ the widely-used five metrics while excluding the self-defined metrics for measuring the fairness of models, i.e., SPD (Mehrabi et al., 2022; Biswas and Rajan, 2020, 2021; Chen et al., 2022b), DEO (Mehrabi et al., 2022; Jung et al., 2021; Zhang et al., 2022b), EOD (Mehrabi et al., 2022; Biswas and Rajan, 2020, 2021; Chen et al., 2022b), AAOD (Biswas and Rajan, 2020, 2021; Chen et al., 2022b), and AED (Biswas and Rajan, 2020, 2021; Mao et al., 2023), following previous work. The reason we chose these five metrics is twofold: (1) They have been integrated into the well-known AIF360 and Fairlearn toolkits, indicating the importance and representativeness of them. (2) They have been well recognized and adopted by at least three previous studies shown in Table  1.
@@
Considering that the classification task for CIFAR-10S involves multi-class classification, whereas the classic fairness metrics mentioned earlier are primarily designed for binary classification tasks, we employ the ”ovr (one-vs-rest)” strategy, inspired by  (Zhang et al., 2022b; Jung et al., 2021) and sklearn to decompose the N-ary classification problem into N binary classification subproblems. In our experiments, we initiate the process by computing the bias score for each class, following the same procedure as described for binary classification. Specifically, we calculate the fairness metrics for each class against the rest and then average the results across different classes to derive the multi-class versions of these fairness metrics.
@@ 4.3. Implementation Details
For all experiments, we use ResNet-18  (He et al., 2016) as the backbone architecture by following existing studies (Qraitem et al., 2022; Sarridis et al., 2023; Jung et al., 2021; Wang et al., 2022; Zhang et al., 2022b). Moreover, we used the best default configurations of the studied techniques provided in the respective papers. To mitigate the effect of randomness. We have also repeated all our experiments 10 times to ensure the reliability and stability of the results, which took about 40 days apart from the costs of tool implementation and configuration. It’s worth noting that since some approaches depend on a well-trained model (i.e., post-processing) while some approaches train the models from scratch (e.g., pre-processing). We compare the performance of the final optimized models directly. For methods that depend on existing well-trained models, we firstly conducted a grid search process to obtain the best-performing models as their inputs.
Both our performance and fairness metrics are implemented based on sklearn (Pedregosa et al., 2011) and Fairlearn (Weerts et al., 2023) libraries. Particularly, we have implemented them in a easy-to-use framework for future studies. It is available in our open-source repository. All experiments are implemented with Python 3.8.17 and PyTorch 1.10.1, and executed on a Ubuntu 18.04.2 LTS with two 2.00GHz Intel(R) Xeon(R) CPU E5-2660 v4 CPUs and four NVIDIA GeForce RTX 2080 Ti GPUs.
@@ 5.1. RQ1: Overall Effectiveness
For further analyzing the reasons of the different effectiveness of methods belong to the same category, we first take the pre-processing category as an example. We believe that the main reason is due to their underlying principles and their suitability for addressing class imbalance and bias in the data. For example, US works well when the class imbalance is significant, as it directly addresses this issue by reducing the majority class. Additionally, by selectively dropping samples from specific subgroups, it can effectively mitigate bias by balancing the sizes of these subgroups. However, OS may result in over-fitting, especially with complex models like deep neural networks, when samples are excessively duplicated. This can lead to poorer generalization on the test set, which might explain its lower performance. To our surprise, as a newly proposed method, BM might not always outperform other techniques, such as US, in terms of fairness. The possible reason is that BM might introduce additional complexity to the pre-processing pipeline by emulating bias distributions across classes. This added complexity may lead to over-fitting or decreased generalization performance. In this case, a simpler method like US might suffice and even outperform BM on fairness. However, as aforementioned, when considering both model fairness and accuracy, BM still should be the best.
@@
Similarly, for in-processing methods, we observe that although we cannot determine which method performs the best, the Adv method exhibits the poorest fairness performance across all datasets, especially on the UTKFace-age as shown in Table 3. The core idea behind Adv is fairness through blindness (Wang et al., 2020), which means attempting to prevent a model from explicitly encoding information about sensitive attributes through adversarial training. However, as we have observed, this practice of ignoring sensitive attribute encoding has several issues. Firstly, due to the phenomenon of redundant encoding (Hardt et al., 2016; Dwork et al., 2012), even if there is no particular sensitive attribute in the classifier’s feature representation, combinations of other attributes can be used as a proxy. For instance, considering a real-world task where a bank assesses a loan application without considering the applicant’s gender. If the applicant’s occupation is nurse, we can still infer with a high probability that the applicant is likely female based on the nurse attribute. Secondly, this practice of discarding sensitive attribute encoding may harm accuracy, and the results in Figure 1 support this hypothesis, with Adv’s lowest accuracy across all datasets among in-processing methods. The reason may be that internal connections between different attributes likely have a significant impact on accuracy.
@@ 5.2. RQ2: Influence of evaluation metrics
Moreover, we conducted a further in-depth investigation about the correlations between different fairness metrics and aimed to provide valuable guidance for future studies. From Figure 2 we can observe that AAOD has the highest overall correlation as its correlation coefficients with the other four metrics never falling below 0.5. We analyze the reasons based on the definitions of each fairness metric (see Section  4.2.1). The goal of SPD is to ensure a DL model’s predictions are independent of membership in a sensitive group. It is an easy metric but it only considers a single binary protected attribute, which is sometimes insufficient. For DEO, EOD and AAOD, the goal of these metrics is to ensure a DL model performs equally well for different groups. These metrics are stricter than SPD because they require consideration of both the sensitive attribute and the true labels, involving the measurement of the difference between true positive rate and false positive rate for the unprivileged and privileged groups. AAOD, as an arithmetic average of the difference in false positive rates and the difference in true positive rates, effectively balances the effects of both difference metrics. This is supported by the experimental results, where AAOD generally exhibits higher differentiation compared to DEO and EOD. Therefore, AAOD appears to be a robust fairness evaluation metric compared to DEO and EOD, especially in scenarios where DEO and EOD show substantial discrepancies. AED measures the difference in misclassification rates of the model on different sensitive attribute groups. The major drawback of this metric is that its values are generally small, particularly on the CIFAR-10S dataset, resulting in lower differentiability compared to SPD and AAOD. Therefore, it is not considered an excellent fairness evaluation metric.
@@
Finding 8: For imbalanced dataset, accuracy and balanced accuracy tend to produce diverse results, and thus both of them should be used to comprehensively measure the performance of fairness improvement methods.
@@ 6. Implications
Combining the strength of different methods. According to our empirical results, there is no best-performing methods in all application scenarios and different methods are conceptually complementary. For example, pre-processing methods focus more on transforming or augmenting training data, while in-processing methods primarily optimize the design of objective functions and model structures. Our experimental results also confirm that. Therefore, it should be promising to combine the strength of individual methods since they improve the models from different perspectives, e.g., training data and training process. In particular, directly combining the results of different methods can be effective. However, investigating the deep combination of different methods by considering their core novelty still needs more exploration.
Developing more effective post-processing methods. With the rapid development and inspiring performance of large language models (LLMs), they have been adopted in various applications. However, re-training LLMs can be a hard or even impossible task for end users due to the high requirement on computing resources and the reliance on training data. Besides, for deployed models, post-processing methods are intuitively the best choice for improving model fairness. Nevertheless, the results shown in Table 3 demonstrate that existing post-processing methods are not satisfactory and still have much room for further improvement compared with the other two categories. Therefore, more studies should focus on the post-processing methods.
Understanding the source of unfairness. Based on the introduction of existing approaches, almost all of them aim to improve the model fairness in somehow a blind way without understanding the core reasons of the unfairness in DL models. Specifically, pre-processing methods assume that the training data contain bias while in-processing methods deem the bias is involved during the training process. However, even adopting the same training data and the same training process, the model performance regarding different attributes can also be different. Therefore, understanding the source of unfairness in the DL models may help to better overcome this issue. In particular, model interpretation techniques (Shen et al., 2021; Chen et al., 2019) can be incorporated to facilitate this task.
@@ 8.2. Empirical Studies of Model Fairness
Empirical studies are essential for evaluating the effectiveness and efficiency of fairness improvement methods. Additionally, it can inspire researchers to pursue further innovations in fairness improvement. Most previous studies (Chakraborty et al., 2019; Friedler et al., 2019; Biswas and Rajan, 2020; Zhang and Sun, 2022) on fairness have focused on numerical or tabular inputs. However, fairness issues in image data are more complex. Wang et al(Wang et al., 2020) conducted an analysis of computer vision models, specifically targeting adversarial training approaches, while Chen et al. (Chen et al., 2024) studied fairness improvement methods that working with tabular inputs. In contrast, our research delves into 13 approaches working with image inputs from three major categories. Additionally, we use three distinct datasets and encompass all commonly-used performance metrics.

### 24 Mar 2024 v2

@@ 4.1. Dataset Selection
As a result, we selected three datasets (i.e., CelebA, UTKFace, and CIFAR-10S) for conducting the empirical study. The summary of the datasets is shown in Table 2, which presents the sensitive attribute considered in the study, the labels of the classification tasks, and the studied methods that previously used the same settings. In addition, we also present the number of instances for model training and testing per each dataset by following existing studies.
@@
Specifically, CelebA and UTKFace datasets represent human-centric datasets that involve sensitive attributes (e.g., age and race) reflected in facial images, while CIFAR-10S is a more general dataset in fairness study, which introduces biases related to image colors of objects (non-facial images). They were previously used for evaluation by partial/all methods under different conditions, e.g., using different metrics. From the table we can also find that previous studies tend to employ different datasets, hindering their comprehensive comparison. In this study, we aim to provide an extensive study through a systematic evaluation by employing all of them, which vary in multiple perspectives.
@@ 4.2.1. Fairness Metrics
Numerous prior studies have predominantly assessed fairness using a restricted set of fairness metrics. Although certain approaches have introduced their unique fairness metrics, these metrics have not achieved widespread adoption and lack broad applicability. To address these limitations, we employ the widely-used five metrics while excluding the self-defined metrics for measuring the fairness of models, i.e., SPD (Mehrabi et al., 2022; Biswas and Rajan, 2020, 2021; Chen et al., 2022b), DEO (Mehrabi et al., 2022; Jung et al., 2021; Zhang et al., 2022), EOD (Mehrabi et al., 2022; Biswas and Rajan, 2020, 2021; Chen et al., 2022b), AAOD (Biswas and Rajan, 2020, 2021; Chen et al., 2022b), and AED (Biswas and Rajan, 2020, 2021; Mao et al., 2023), following previous work. The reason we chose these five metrics is twofold: (1) They have been integrated into the well-known AIF360 and Fairlearn toolkits, indicating the importance and representativeness of them. (2) They have been well recognized and adopted by at least three previous studies shown in Table  1.
@@
Considering that the classification task for CIFAR-10S involves multi-class classification, whereas the classic fairness metrics mentioned earlier are primarily designed for binary classification tasks, we employ the ”ovr (one-vs-rest)” strategy, inspired by  (Zhang et al., 2022; Jung et al., 2021) and sklearn to decompose the N-ary classification problem into N binary classification subproblems. In our experiments, we initiate the process by computing the bias score for each class, following the same procedure as described for binary classification. Specifically, we calculate the fairness metrics for each class against the rest and then average the results across different classes to derive the multi-class versions of these fairness metrics.
@@ 4.3. Implementation Details
For all experiments, we use ResNet-18  (He et al., 2016) as the backbone architecture by following existing studies (Qraitem et al., 2022; Sarridis et al., 2023; Jung et al., 2021; Wang et al., 2022; Zhang et al., 2022). Moreover, we used the best default configurations of the studied techniques provided in the respective papers. Specifically, we firstly preprocessed the selected three datasets to ensure uniform data partitioning for consistency. Then, for each studied method, we utilized the optimal default configuration hyper-parameters provided in their respective papers to ensure the reliability of the experiments. However, since some methods originally did not utilize all of the studied datasets used in this paper, we adhered to their principles and adapted them to the new datasets , fine-tuning them for optimal results. In summary, to effectively and fairly compare the experimental results of each method, we standardized various factors such as data partitioning, model backbone, experiment repetition, experimental environment and so on. In particular, to mitigate the effect of randomness, we have also repeated all our experiments 10 times to ensure the reliability and stability of the results, which took about 40 days apart from the costs of tool implementation and configuration. It’s worth noting that since some approaches depend on a well-trained model (i.e., post-processing) while some approaches train the models from scratch (e.g., pre-processing), we compare the performance of the final optimized models directly. For methods that depend on existing well-trained models, we first conducted a grid search process to obtain the best-performing models as their inputs. We have published all the implementation details in our open-source repository.
Both our performance and fairness metrics are implemented based on sklearn (Pedregosa et al., 2011) and Fairlearn (Weerts et al., 2023) libraries. Particularly, we have implemented them in an easy-to-use framework for future studies. It is available in our open-source repository. All experiments are implemented with Python 3.8.18 and PyTorch 1.10.1, and executed on an Ubuntu 20.04.6 LTS with 2.90GHz Intel(R) Xeon(R) Gold 6326 CPUs and eight NVIDIA GeForce RTX 3090 GPUs.
@@ 5.1. RQ1: Overall Effectiveness
For further analysis of the reasons for the different effectiveness of methods belonging to the same category, we first take the pre-processing category as an example. We believe that the main reason is due to their underlying principles and their suitability for addressing class imbalance and bias in the data. For example, US works well when the class imbalance is significant, as it directly addresses this issue by reducing the majority class. Additionally, by selectively dropping samples from specific subgroups, it can effectively mitigate bias by balancing the sizes of these subgroups. However, OS may result in over-fitting, especially with complex models like deep neural networks, when samples are excessively duplicated. This can lead to poorer generalization on the test set, which might explain its lower performance. To our surprise, as a newly proposed method, BM might not always outperform other techniques, such as US, in terms of fairness. The possible reason is that BM might introduce additional complexity to the pre-processing pipeline by emulating bias distributions across classes. This added complexity may lead to over-fitting or decreased generalization performance. In this case, a simpler method like US might suffice and even outperform BM on fairness. However, as aforementioned, when considering both model fairness and accuracy, BM should still be the best.
@@
Similarly, for in-processing methods, we observe that although we cannot determine which method performs the best, the Adv method exhibits the poorest fairness performance across all datasets, especially on the UTKFace-age as shown in Table 3. The core idea behind Adv is fairness through blindness (Wang et al., 2020a), which means attempting to prevent a model from explicitly encoding information about sensitive attributes through adversarial training. However, as we have observed, this practice of ignoring sensitive attribute encoding has several issues. Firstly, due to the phenomenon of redundant encoding (Hardt et al., 2016; Dwork et al., 2012), even if there is no particular sensitive attribute in the classifier’s feature representation, combinations of other attributes can be used as a proxy. For instance, considering a real-world task where a bank assesses a loan application without considering the applicant’s gender. If the applicant’s occupation is nurse, we can still infer with a high probability that the applicant is likely female based on the nurse attribute. Secondly, this practice of discarding sensitive attribute encoding may harm accuracy, and the results in Figure 1 support this hypothesis, with Adv’s lowest accuracy across all datasets among in-processing methods. The reason may be that internal connections between different attributes likely have a significant impact on accuracy.
@@ 5.2. RQ2: Influence of evaluation metrics
Moreover, we conducted a further in-depth investigation into the correlations between different fairness metrics and aimed to provide valuable guidance for future studies. From Figure 2 we can observe that AAOD has the highest overall correlation as its correlation coefficients with the other four metrics never fall below 0.5. We analyze the reasons based on the definitions of each fairness metric (see Section  4.2.1). The goal of SPD is to ensure a DL model’s predictions are independent of membership in a sensitive group. It is an easy metric but it only considers a single binary protected attribute, which is sometimes insufficient. For DEO, EOD and AAOD, the goal of these metrics is to ensure a DL model performs equally well for different groups. These metrics are stricter than SPD because they require consideration of both the sensitive attribute and the true labels, involving the measurement of the difference between the true positive rate and false positive rate for the unprivileged and privileged groups. AAOD, as an arithmetic average of the difference in false positive rates and the difference in true positive rates, effectively balances the effects of both difference metrics. This is supported by the experimental results, where AAOD generally exhibits higher differentiation compared to DEO and EOD. Therefore, AAOD appears to be a robust fairness evaluation metric compared to DEO and EOD, especially in scenarios where DEO and EOD show substantial discrepancies. AED measures the difference in misclassification rates of the model on different sensitive attribute groups. The major drawback of this metric is that its values are generally small, particularly on the CIFAR-10S dataset, resulting in lower differentiability compared to SPD and AAOD. Therefore, it is not considered an excellent fairness evaluation metric.
@@
Finding 8: For an imbalanced dataset, accuracy and balanced accuracy tend to produce diverse results, and thus both of them should be used to comprehensively measure the performance of fairness improvement methods.
@@ 6. Implications
Combining the strength of different methods. According to our empirical results (i.e. Finding 1), there is no best-performing methods in all application scenarios and different methods are conceptually complementary. For example, pre-processing methods focus more on transforming or augmenting training data, while in-processing methods primarily optimize the design of objective functions and model structures. Our experimental results also confirm that. Therefore, it should be promising to combine the strength of individual methods since they improve the models from different perspectives, e.g., training data and training process. In particular, directly combining the results of different methods can be effective. However, investigating the deep combination of different methods by considering their core novelty still needs more exploration.
Developing more effective post-processing methods. With the rapid development and inspiring performance of large language models (LLMs), they have been adopted in various applications. However, re-training LLMs can be a hard or even impossible task for end users due to the high requirement on computing resources and the reliance on training data. Besides, for deployed models, post-processing methods are intuitively the best choice for improving model fairness. Nevertheless, according to Finding 5, the results shown in Table 3 demonstrate that existing post-processing methods are not satisfactory and still have much room for further improvement compared with the other two categories. Therefore, more studies should focus on the post-processing methods.
Understanding the source of unfairness. Based on the introduction of existing approaches, almost all of them aim to improve the model fairness in somehow a blind way without understanding the core reasons of the unfairness in DL models. Specifically, pre-processing methods assume that the training data contain bias while in-processing methods deem the bias is involved during the training process. However, even adopting the same training data and the same training process (see Section 5.3), the model performance regarding different attributes can also be different. Therefore, understanding the source of unfairness in the DL models may help to better overcome this issue. In particular, model interpretation techniques (Shen et al., 2021; Chen et al., 2019) can be incorporated to facilitate this task.
Incorporating the correlations among different attributes deeply. Indeed, sensitive attributes play a crucial role in model predictions regarding result fairness. However, according to Finding 4, simply removing the sensitive attributes does not necessarily improve model fairness, and incorporating the correlations between the sensitive attributes and the others is also important. In fact, the existing study (Li et al., 2022) has already considered such information for improving model fairness. Nevertheless, it was designed for the structured (i.e., tabular) data, which should be much simpler than the images since the image features are not clearly defined. In other words, how to effectively identify the implicit correlations between different image features and incorporate them into the process of fairness improvement still requires more in-depth exploration.
@@ 8.2. Empirical Studies of Model Fairness
Empirical studies are essential for evaluating the effectiveness and efficiency of fairness improvement methods. Additionally, it can inspire researchers to pursue further innovations in fairness improvement. Most previous studies (Chakraborty et al., 2019; Friedler et al., 2019; Biswas and Rajan, 2020; Zhang and Sun, 2022) on fairness have focused on numerical or tabular inputs. However, fairness issues in image data are more complex. Wang et al(Wang et al., 2020a) conducted an analysis of computer vision models, specifically targeting adversarial training approaches, while Chen et al. (Chen et al., 2024) studied fairness improvement methods that work with tabular inputs. In contrast, our research delves into 13 approaches working with image inputs from three major categories. Additionally, we use three distinct datasets and encompass all commonly-used performance metrics.
@@ 9. Conclusion
Data Availability
For the sake of Open Science, we make the replication package with source code and the experimental results publicly available at: https://github.com/junjie1003/DL-Fairness-Study.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant Nos. 62202324 and 62322208, and CCF Young Elite Scientists Sponsorship Program (by CAST).